OpenStudy (anonymous):

If f(x) = x^3 - x - 6 for all real numbers x, and if g is the inverse function of f, then f'(g(0))g'(0) = ? a) 0 b) 1 c) -1 d) -6 e) (6)^(1/3) I had a lot of trouble with finding the inverse, If anyone can help me that would be awesome!

OpenStudy (anonymous):

I don't think you need to find the inverse. If f(x) and g(x) are inverses, then \[g(x)=\frac{ 1 }{ f'(g(x)) }\] Multiply both sides by \(f'(g(x))\) and sub in 0 for x

OpenStudy (blacksteel):

You actually don't need to find the inverse function to solve this. It uses the fact that \[[f^{-1}]'(x) = \frac{ 1 }{ f'(f^{-1}(x))}\]In this case, this means that \[g'(x) = \frac{ 1 }{ f'(g(x)) }\]So then \[f'(g(0))g'(0) = \frac{ f'(g(0)) }{ f'(g(0)) } = 1\]

OpenStudy (anonymous):

sorry , should read g'(x) =...